# Uniform Loss vs. Specialized Optimization: A Comparative Analysis in Multi-Task Learning

### v1

@@ Abstract
Specialized Multi-Task Optimizers (SMTOs) balance task learning in Multi-Task Learning by addressing issues like conflicting gradients and differing gradient norms, which hinder equal-weighted task training. However, recent critiques suggest that equally weighted tasks can achieve competitive results compared to SMTOs, arguing that previous SMTO results were influenced by poor hyperparameter optimization and lack of regularization. In this work, we evaluate these claims through an extensive empirical evaluation of SMTOs, including some of the latest methods, on more complex multi-task problems to clarify this behavior. Our findings indicate that SMTOs perform well compared to uniform loss and that fixed weights can achieve competitive performance compared to SMTOs. Furthermore, we demonstrate why uniform loss perform similarly to SMTOs in some instances. The code will be made publicly available.
@@ 1 Introduction
  • An investigation into why equally weighted tasks can match the results of SMTOs. More specifically, the two similar tasks scenario in Section  3.2, and a complexity analysis in Section  3.3.
@@ 3 Experimental Evaluation
where $l_{i,j}$ is $1$ if a lower value is better and 00 otherwise. Likewise, the mean rank is given by:
@@ 3.3 Complexity analysis
[…]
@@ 3.4 QM9
We use the same configuration from [10]. The example from PyTorch Geometric [43] was adapted, and we used the model from [44]. The model was used to predict 11 properties from the molecules. We use the commonly adopted distribution of 110K molecules for training, 10K for validation, and 10K for testing, with 120 of batch size, the learning rate was optimized in $lr$in $5\times 10^{-3}$, $10^{-3}$, $5\times 10^{-4}$}, and the ground truth values were normalized.
@@ A.3 Rotograd’s instability on the QM9 dataset
Figure 9: Mean rotation loss $\left(\mathcal{L}_{rot}\right)$ from RotoGrad. The $\mathcal{L}_{rot}$ is at least 10$\times$ higher than the other cases, which could explain the training instability.
@@ A.4 Proofs
[…]
@@
[…]
@@
[…]
@@
$d^{*}=\left(\frac{1}{A}+\frac{1}{A}\right)^{-1}\left(\frac{g_{1}}{A}+\frac{g_{2}}{A}\right)=\frac{1}{2}(g_{1}+g_{2})\$
@@
$w_{t}=w_{i,t}=c_{t}\frac{z_{i,t}}{\ell_{i,t}},\quad\text{where}\quad c_{t}=\left(\sum_{i}^{N}\frac{z_{i,t}}{\ell_{i},t}\right),\quad\text{for all }i=1,\dots,N\text{ and for all }t.$
@@
$[\delta_{t}]_{i}=D\quad\text{for all }i.$
@@ A.6 Hyperparameters
For the hyperparameters that were optimized without a defined range, such as Auto-Lambda’s and RotoGrad’s auxiliary learning rate, we opted to define it as a scale of the learning rate ($aux\_lr=lr*lr\_scale$).
@@
Table 7: Hyperparameters for MNIST - CL & RR.
@@
Table 8: Hyperparameters for MNIST - RL & RR.
@@
City. S D I Single Task ACC: 0.001 mIoU: 0.001 L1 abs: 0.001 L1 rel: 0.0005 L1 abs: 0.001 0 0 0 0 0 $41e-$ 4 $41e-$ 4 $41e-$ 4 $41e-$ 4 $51e-$ 5
Unit. Scal. 0.001 0 $41e-$ 4
Auto-Lambda 0.001 0 $41e-$ 4 auxiliary learning rate scale: {100 10 1 0.1 0.01 0.001 0.0001} initial weight: 0.1
CAGrad 0.001 0 $41e-$ 4 $c$: {0.2, 0.5, 0.8}
CDTT 0.001 0 $41e-$ 4 $\alpha$: {0.2, 0.4, 0.6, 0.8, 1.0}
EDM 0.001 0 $41e-$ 4
FAMO 0.001 0 $41e-$ 4 $\gamma$: {0.01, 0.001, 0.0001} learning rate of the task logits: 0.025
IMTL 0.001 0 $41e-$ 4
Nash-MTL 0.001 0 $41e-$ 4 optimizer iterations: 20
@@
QM9 $C_{v}$ $G$ $H$ $U_{0}$ $U$ $\alpha$ $\epsilon_{homo}$ $\epsilon_{lumo}$ $\mu$ $R^{2}$ ZPVE Single Task 0.001 0.001 0.001 0.0005 0.001 0.005 0.001 0.001 0.001 0.001 0.001 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0

### v2

@@ Abstract
Specialized Multi-Task Optimizers (SMTOs) balance task learning in Multi-Task Learning by addressing issues like conflicting gradients and differing gradient norms, which hinder equal-weighted task training. However, recent critiques suggest that equally weighted tasks can achieve competitive results compared to SMTOs, arguing that previous SMTO results were influenced by poor hyperparameter optimization and lack of regularization. In this work, we evaluate these claims through an extensive empirical evaluation of SMTOs, including some of the latest methods, on more complex multi-task problems to clarify this behavior. Our findings indicate that SMTOs perform well compared to uniform loss and that fixed weights can achieve competitive performance compared to SMTOs. Furthermore, we demonstrate why uniform loss perform similarly to SMTOs in some instances. The source code is available at https://github.com/Gabriel-SGama/UnitScal_vs_SMTOs.
@@ 1 Introduction
  • An investigation into why equally weighted tasks can match the results of SMTOs. More specifically, the two similar tasks scenario in Section  3.2, and a complexity analysis in Section  $3.$
@@ 3 Experimental Evaluation
where $l_{i,j}$ is $1$ if a lower value is better and 0 otherwise. Likewise, the mean rank is given by:
@@ 3.3 Complexity analysis
We vary model size, number of tasks, and information available to learn to identify which characteristic is relevant. We first trained the model on the same initial two-task configuration of [24] and applied small cumulative steps until the most complex configuration. We opted to test only IMTL [11] and compare it to the Unit. Scal. in the transitional steps as it was one of the best performing SMTOs, and it does not require any specific hyperparameter optimization. All Cityscapes configurations are optimized varying the learning rate at lrlr $\in$ {0.005, 0.001, 0.0005, 0.0001} and the weight decay $\in$ $10^{-4}$, $10^{-5}$, 0.0} with 32 batch size. Each configuration is trained once, and then the best in the validation dataset is trained a total of 3 times. The selected steps are:
@@ 3.4 QM9
We use the same configuration from [10]. The example from PyTorch Geometric [43] was adapted, and we used the model from [44]. The model was used to predict 11 properties from the molecules. We use the commonly adopted distribution of 110K molecules for training, 10K for validation, and 10K for testing, with 120 of batch size, the learning rate was optimized in lrlr $\in$ $5\times 10^{-3}$, $10^{-3}$, $5\times 10^{-4}$}, and the ground truth values were normalized.
@@ A.3 Rotograd’s instability on the QM9 dataset
[…]
@@ A.4 Proofs
$\left\{\begin{aligned} \alpha_{1}g_{1}^{\top}g_{1}+\alpha_{2}g_{2}^{\top}g_{1}&=\frac{1}{\alpha_{1}}\\ \alpha_{2}g_{2}^{\top}g_{2}+\alpha_{1}g_{1}^{\top}g_{2}&=\frac{1}{\alpha_{2}}\end{aligned}\right$.
@@
$\left\{\begin{aligned} \alpha_{1}A^{2}+\alpha_{2}A^{2}p&=\frac{1}{\alpha_{1}}\\ \alpha_{2}A^{2}+\alpha_{1}A^{2}p&=\frac{1}{\alpha_{2}}\end{aligned}\right$.
@@
$\left\{\begin{aligned} \alpha_{1}^{2}A^{2}+\alpha_{1}\alpha_{2}A^{2}p&=1\\ \alpha_{2}^{2}A^{2}+\alpha_{1}\alpha_{2}A^{2}p&=1\end{aligned}\right$.
@@
$d^{*}=\left(\frac{1}{A}+\frac{1}{A}\right)^{-1}\left(\frac{g_{1}}{A}+\frac{g_{2}}{A}\right)=\frac{1}{2}(g_{1}+g_{2})\\$
@@
[…]
@@
$\delta_{t}]_{i}=D\quad\text{for all }i.$
@@ A.6 Hyperparameters
For the hyperparameters that were optimized without a defined range, such as Auto-Lambda’s and RotoGrad’s auxiliary learning rate, we opted to define it as a scale of the learning rate (aux_lr=lrlr_scaleaux\_lr=lr*lr\_scale).
@@
Table 7: Hyperparameters for MNIST - CL & $R$.
@@
Table 8: Hyperparameters for MNIST - RL & $R$.
@@
City. S D I Single Task ACC: 0.001 mIoU: 0.001 L1 abs: 0.001 L1 rel: 0.0005 L1 abs: 0.001 0 0 0 0 0 1e41e-4 1e41e-4 1e41e-4 1e41e-4 1e51e-5
Unit. Scal. 0.001 0 1e41e-4
Auto-Lambda 0.001 0 1e41e-4 auxiliary learning rate scale: {100 10 1 0.1 0.01 0.001 0.0001} initial weight: 0.1
CAGrad 0.001 0 1e41e-4 $c$: {0.2, 0.5, 0.8}
CDTT 0.001 0 1e41e-4 $\alpha$: {0.2, 0.4, 0.6, 0.8, 1.0}
EDM 0.001 0 1e41e-4
FAMO 0.001 0 1e41e-4 $\gamma$: {0.01, 0.001, 0.0001} learning rate of the task logits: 0.025
IMTL 0.001 0 1e41e-4
Nash-MTL 0.001 0 1e41e-4 optimizer iterations: 20
@@
QM9 $C_{v}$ $G$ $H$ $_{0} UU$ $\alpha$ $\epsilon_{homo}$ $\epsilon_{lumo}$ $\mu$ $R^{2}$ ZPVE Single Task 0.001 0.001 0.001 0.0005 0.001 0.005 0.001 0.001 0.001 0.001 0.001 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0